Bibliography Hans–Paul Schwefel

August 3, 2022

[1] Hans-Paul Schwefel. Kybernetische Evolution als Strategie der experimentellen Forschung in der Strömungstechnik. Dipl.-Ing. Thesis, Technical University of Berlin, Hermann Föttinger–Institute for Fluid Dynamics, March 1965.

[2] Peter Bienert, Ingo Rechenberg, and Hans-Paul Schwefel. Messung kleiner Wandschubspannungen bei turbulenten Grenzschichten in Ablösenähe. Technical Report of the Hermann Föttinger– Institute for Fluid Dynamics Wi 8/45, Technical University of Berlin, September 1966.

[3] Reinhard Radebold, Hans-Paul Schwefel, et al. Zum Stand der eigenen und ausländischen Arbeiten (Summer 1968, Warschau-Symposium), part 1. Technical Report No. 30 of the Project MHD– Staustrahlrohr 11.028/68, AEG Research Institute, Berlin, September 1968.

[4] Hans-Paul Schwefel. Experimentelle Optimierung einer Zweiphasendüse, part I. Technical Report No. 35 of the Project MHD–Staustrahlrohr 11.034/68, AEG Research Institute, Berlin, October 1968.

[5] Hans-Paul Schwefel. Großer MHD–Versuchsstand, Na–Erhitzer. Technical Report No. 42 of the Project MHD–Staustrahlrohr 11.041/68, AEG Research Institute, Berlin, October 1968.

[6] Hans-Paul Schwefel and Dietrich Weinberg. Untersuchungen über gekühlte Zweiphasendüsen, part I. Technical Report No. 32 of the Project MHD–Staustrahlrohr 11.031/68, AEG Research Institute, Berlin, September 1968.

[7] Jürgen Klockgether and Hans-Paul Schwefel. Two–phase nozzle and hollow core jet experiments. In D. G. Elliott, editor, Proc. Eleventh Symp. Engineering Aspects of Magnetohydrodynamics, pages 141–148, California Institute of Technology, Pasadena CA, 1970.

[8] Hans-Paul Schwefel. Wärmeübertragungsanlage für Flüssigmetall bis 900^{∘}C. VDI–Berichte No. 153, Wärmeübertragungs-Anlagen, pages 47–52, 1970.

[9] Hans-Paul Schwefel. Untersuchungen über gekühlte Zweiphasendüsen, part III. Technical Report No. 32 of the Project MHD–Staustrahlrohr 11.002/70, AEG Research Institute, Berlin, January 1970.

[10] Hans-Paul Schwefel. Evolutionsstrategie für die numerische Optimierung (1). Technical Report of the Institute of Measurement and Control Technology and of the Institute of Biology and Anthropology Ga 24/16, Technical University of Berlin, November 1970.

[11] Reinhard Radebold, Hans-Paul Schwefel, et al. Energy conversion with liquid metal working fluids in the MHD–Staustrahlrohr. In Proc. Fifth Int’l Conf. Magnetohydrodynamic Electrical Power Generation, Munich, volume 3, pages 195–214, IAEA, Vienna, 1971.

[12] Hans-Paul Schwefel. Evolutionsstrategie für die numerische Optimierung (2). Technical Report of the Institute of Measurement and Control Technology and of the Institute of Biology and Anthropology Ga 24/16, Technical University of Berlin, May 1971.

[13] Hans-Paul Schwefel. Evolutionsstrategie für die numerische Optimierung (3). Technical Report of the Institute of Measurement and Control Technology and of the Institute of Biology and Anthropology Ga 24/21, Technical University of Berlin, May 1974.

[14] Hans-Paul Schwefel. Adaptive Mechanismen in der biologischen Evolution und ihr Einfluss auf die Evolutionsgeschwindigkeit. Technical Report of the Working Group of Bionics and Evolution Techniques at the Institute for Measurement and Control Technology Re 215/3, Technical University of Berlin, July 1974.

[15] Hans-Paul Schwefel. Evolutionsstrategie und numerische Optimierung. Dr.-Ing. Thesis, Technical University of Berlin, Department of Process Engineering, 1975.

[16] Hans-Paul Schwefel. Binäre Optimierung durch somatische Mutation. Technical Report of the Working Group of Bionics and Evolution Techniques at the Institute of Measurement and Control Technology of the Technical University of Berlin and of the Central Animal Laboratory of the Medical Highschool of Hanover, SFB 146 Versuchstierforschung at the Veterinary Highschool of Hanover, May 1975.

[17] Kurt Schmitz, Ulrich Schöler, and Hans-Paul Schwefel. Dynamische Energiemodelle – eine brauchbare Entscheidungshilfe? Energiewirtschaftliche Tagesfragen, 28:728–729, 1976.

[18] Kurt Schmitz and Hans-Paul Schwefel. Finding reasonable energy policies by means of a dynamic simulation model. In M. H. Hamza, editor, Proc. Int’l Symp. Simulation’77, Montreux, pages 248–253, Acta Press, Anaheim CA, 1977.

[19] Kurt Schmitz and Hans-Paul Schwefel. An integrated simulation model of the Federal Republic of Germany as a decision aid for analyzing and planning the energy system. In H. J. Highland, R. G. Sargent, and J. W. Schmidt, editors, Proc. Winter Simulation Conf., volume 1, pages 202–211, Gaithersburg MD, 1977.

[20] Hans-Paul Schwefel. Numerische Optimierung von Computer–Modellen mittels der Evolutionsstrategie, volume 26 of Interdisciplinary Systems Research. Birkhäuser, Basle, 1977.

[21] Rainer Heckler and Hans-Paul Schwefel. Superimposing direct search methods for parameter optimization onto dynamic simulation models. In H. J. Highland, N. R. Nielsen, and L. G. Hull, editors, Proc. Winter Simulation Conf., volume 1, pages 173–181, Miami Beach FL, 1978.

[22] Hans-Paul Schwefel. Optimierung von Simulationsmodellen mit der Evolutionsstrategie. In B. Schneider and U. Ranft, editors, Simulationsmethoden in der Medizin und Biologie, pages 115– 129. Springer, Berlin, 1978.

[23] Friedhelm Drepper, Rainer Heckler, and Hans-Paul Schwefel. Ein integriertes System von Schätzverfahren, Simulations- und Optimierungstechniken zur rechnergestützten Langfristplanung. In K. H. Böhling and P. P. Spies, editors, German Society for Informatics (GI) – 9. Jahrestagung, pages 296–308, Springer, Berlin, 1979.

[24] Hans-Paul Schwefel. Direct search for optimal parameters within simulation models. In R. D. Conine, E. D. Katz, and J. E. Melde, editors, Proc. Twelfth Annual Simulation Symp., Tampa FL, pages 91–102, IEEE Computer Society, Long Beach CA, 1979.

[25] Hans-Paul Schwefel. Ein Leistungsvergleich ableitungsfreier Methoden der nichtlinearen Parameteroptimierung. In J. Schwarze, L. von Dobschütz, et al., editors, Proc. Operations Research 9, DGOR Jahrestagung 1979, pages 503–504, Physica–Verlag, Würzburg, 1980.

[26] Hans-Paul Schwefel. Optimierende Simulation – ein Instrumentarium für die Langfristplanung. In A. Voß and K. Schmitz, editors, Energiemodelle für die Bundesrepublik Deutschland, pages 221–241. TÜV Rheinland, Cologne, 1980.

[27] Hans-Paul Schwefel. Subroutines EVOL, GRUP, KORR – Listings and user’s guides. Technical Report of the Programme Group of Systems Analysis and Technological Development KFA-STE-IB-2/80, Nuclear Research Centre (KFA) Jülich, April 1980.

[28] Hans-Paul Schwefel. Unterprogramme EVOL, GRUP, KORR – Programme und Benutzeranleitungen. Technical Report of the Programme Group of Systems Analysis and Technological Development KFA-STE-IB-3/80, Nuclear Research Centre (KFA) Jülich, April 1980.

[29] Friedhelm Drepper, Rainer Heckler, and Hans-Paul Schwefel. Energieplanung mittels dynamischer Simulation und überlagerter Optimierung. In D. Fischer, H.-Chr. Pfohl, K.-P. Schuster, and J. Schwarze, editors, Proc. Operations Research, DGOR Jahrestagung 1980, pages 545–547, Springer, Berlin, 1981.

[30] Friedhelm Drepper, Rainer Heckler, and Hans-Paul Schwefel. Investment policy – A multicriteria optimal response to present and future oil prices. In J. P. Brans, editor, Operational Research’81, Ninth IFORS Conf., Hamburg, volume Abstracts, pages A102–A103, North–Holland, Amsterdam, 1981.

[31] Gerd Egberts, Rainer Heckler, Hans-Paul Schwefel, and Alfred Voß. Integration of optimization and simulation models. In I. Kavrakoglu, editor, Mathematical Modelling of Energy Systems, pages 189–207. Sijthoff and Noordhoff, Alphen aan den Rijn, The Netherlands, 1981.

[32] Kurt Schmitz and Hans-Paul Schwefel. Energy system management by computer aided analysis and synthesis. In W. Häfele and L. K. Kirchmeyer, editors, Modeling of large–scale energy systems, pages 223–229. Pergamon Press, Oxford, 1981.

[33] Hans-Paul Schwefel. Numerical Optimization of Computer Models. Wiley, Chichester, 1981.

[34] Hans-Paul Schwefel. Optimum-Suchverfahren – Benutzeranleitung. Technical Report of the Programme Group of Systems Analysis and Technological Development KFA-STE-IB-1/81, Nuclear Research Centre (KFA) Jülich, January 1981.

[35] Hans-Paul Schwefel. Optimum seeking methods – User’s guides. Technical Report of the Programme Group of Systems Analysis and Technological Development KFA-STE-IB-7/81, Nuclear Research Centre (KFA) Jülich, October 1981.

[36] Hans-Paul Schwefel, Rainer Heckler, and Friedhelm Drepper. Combining estimation, simulation, and optimization in computer-aided energy planning. Angewandte Systemanalyse, 2:67–79, 1981.

[37] Hans-Paul Schwefel, Barbara Röth, Margot Köhnen, and Vinzenz Bundschuh. EBAS – Energiebilanz-Auskunftssystem: Anleitung zur Benutzung der Energiebilanzen der Bundesrepublik Deutschland seit 1950. Technical Report of the Programme Group of Systems Analysis and Technological Development KFA-STE-IB-6/81, Nuclear Research Centre (KFA) Jülich, October 1981.

[38] Eckhard Höpfinger and Hans-Paul Schwefel. Trade between industrialized and resource-rich countries – A game-theoretic approach. In W. Eichhorn, R. Henn, K. Neumann, and R. W. Shephard, editors, Economic Theory of Natural Resources, pages 175–185. Physica–Verlag, Würzburg, 1982.

[39] Hans-Paul Schwefel. Evolution und Optimierung. KFA Jahresbericht, 1981/82:67–77, 1982. Nuclear Research Centre (KFA) Jülich.

[40] Friedhelm Drepper and Hans-Paul Schwefel. Ein einfaches dynamisches Investitionsverflechtungsmodell mit nichtlinearen Eigenschaften. In G. Eilenberger and H. Müller-Krumbhaar, editors, Summer Course ’83 ‘Nichtlineare Dynamik in kondensierter Materie’, chapter XXX, pages 1–17. Nuclear Research Centre (KFA) Jülich, 1983.

[41] Helmut Kollmann, Hans-Paul Schwefel, et al. Wie funktioniert das? Die Energie (a dictionary). Bibliographisches Institut, Mannheim, 1983.

[42] Hans-Paul Schwefel. Aktualisierung des BMWi-Energieumwandlungsmodells – Anleitung und Dokumentation für den Zeitraum 1977–1981. Technical Report of the Programme Group of Systems Analysis and Technological Development KFA-STE-IB-1/83, Nuclear Research Centre (KFA) Jülich, December 1983.

[43] Hans-Paul Schwefel. Evolution strategies: A family of non-linear optimization techniques based on imitating some principles of organic evolution. Annals of Operations Research, 1:165–167, 1984.

[44] Friedhelm Drepper and Hans-Paul Schwefel. Ein einfaches dynamisches Investitionsverflechtungsmodell mit nichtlinearen Eigenschaften. In G. Eilenberger and H. Müller-Krumbhaar, editors, Summer Course ’83 ‘Nichtlineare Dynamik in kondensierter Materie’, chapter XXX, pages 1–17. Nuclear Research Centre (KFA) Jülich, 2nd edition, 1985.

[45] Hans-Paul Schwefel. Collective phenomena in evolutionary systems. In P. Checkland and I. Kiss, editors, Problems of Constancy and Change – The Complementarity of Systems Approaches to Complexity, Proc. 31st Annual Meeting, volume 2, pages 1025–1033, Int’l Soc. for General System Research, Budapest, 1987.

[46] Hans-Paul Schwefel. Collective intelligence in evolving systems. In W. Wolff, C. J. Soeder, and F. Drepper, editors, Ecodynamics, Contributions to Theoretical Ecology, pages 95–100. Springer, Berlin, 1988.

[47] Hans-Paul Schwefel. Optimum seeking by imitating natural intelligence. In R. Vichnevetsky, P. Borne, and J. Vignes, editors, Proc. Twelfth IMACS World Congress Scientific Computation, volume 2, pages 52–55, Int’l Assoc. Mathematics and Computers in Simulation (IMACS), Paris, 1988.

[48] Hans-Paul Schwefel. Towards large–scale long–term systems analysis. In W.-M. Cheng, editor, Proc. Int’l Conf. Systems Science and Engineering (ICSSE’88), Beijing, pages 375–381, Int’l Academic Publ. / Pergamon Press, Oxford, 1988.

[49] Hans-Paul Schwefel. Evolutionary learning optimum–seeking on parallel computer architectures. In A. Sydow, S. G. Tzafestas, and R. Vichnevetsky, editors, Systems Analysis and Simulation, volume 1, pages 217–225. Akademie–Verlag, Berlin, 1988.

[50] Ivan Campos, Ernst Peters, and Hans-Paul Schwefel. Zwei Beiträge zum wissensbasierten Einsatz von Optimumsuchverfahren. Technical Report ‘Grüne Reihe’ No. 311, University of Dortmund, Department of Computer Science, 1989.

[51] Ivan Campos and Hans-Paul Schwefel. KBOPT: A knowledge based optimization system. In C. A. Brebbia and S. Hernandez, editors, Computer Aided Optimum Design of Structures: Applications, pages 211–221. Springer, Berlin, 1989.

[52] Hans-Paul Schwefel. Natürliche Intelligenz in evolutionären Systemen. In J. Albertz, editor, Evolution und Evolutionsstrategien in Biologie, Technik und Gesellschaft, volume 9, pages 151–164. Schriftenreihe der Freien Akademie, Wiesbaden, 1989.

[53] Hans-Paul Schwefel. Simulation evolutionärer Lernprozesse. In D. P. F. Möller, editor, Erwin– Riesch–Workshop on Systems Analysis of Biomedical Processes, Bad Münster am Stein, 1988, pages 17–30, Vieweg, Braunschweig, 1989.

[54] Hans-Paul Schwefel. Collective learning by artificial evolution. In P. Bock, F. J. Radermacher, and M. M. Richter, editors, Proc. FAW Workshop Adaptive Learning, Schloss Reisensburg, pages 177–182, Research Institute for Applied Knowledge Engineering (FAW), Ulm, 1989.

[55] Hans-Paul Schwefel. Technische versus biologische Optimierung – Kann man aus der Analogie etwas lernen? In R. Reiner and H. Wirth, editors, Proc. First Int’l Symp. SFB 230 Natürliche Konstruktionen – Leichtbau in Architektur und Natur, 1988, volume 2, pages 155–159, Universities of Stuttgart and Tübingen, Stuttgart, 1989.

[56] Thomas Bäck, Frank Hoffmeister, Frank Kursawe, Günter Rudolph, and Hans-Paul Schwefel. Four contributions to the development of evolution strategies. Technical Report ‘Grüne Reihe’ No. 368, University of Dortmund, Department of Computer Science, December 1990.

[57] Ulrich Hermes, Frank Hoffmeister, Axel Leucht, Hans-Paul Schwefel, and Günter Schwichtenberg. Ein Konzept zum Aufbau eines Campus–weiten Transputer-Systems. Technical Report of the Computing Center (HRZ), University of Dortmund, May 1990.

[58] Frank Hoffmeister and Hans-Paul Schwefel. A taxonomy of parallel evolutionary algorithms. In G. Wolf, T. Legendi, and U. Schendel, editors, Parcella’90, Proc. Fifth Int’l Workshop Parallel Processing by Cellular Automata and Arrays, pages 97–107, Akademie–Verlag, Berlin, 1990.

[59] Frank Hoffmeister and Hans-Paul Schwefel. KORR 2.1 – An implementation of a (μ ,^{+} λ)-evolution strategy. Technical Report, University of Dortmund, Department of Computer Science, November 1990.

[60] Ernst Peters and Hans-Paul Schwefel. Technische Optimierung auf Parallelrechnern. Neue Systeme – Technik für Menschen: Technologieregion Ruhrgebiet, Systec:44/85, 1990.

[61] Hans-Paul Schwefel. Systems analysis, systems design, and evolutionary strategies. Systems Analysis – Modelling – Simulation, 7:853–864, 1990.

[62] Hans-Paul Schwefel. Some observations about evolutionary optimization algorithms. In H.-M. Voigt, H. Mühlenbein, and H.-P. Schwefel, editors, Evolution and Optimization ’89 – Selected Papers on Evolution Theory, Combinatorial Optimization, and Related Topics, Wartburg, 1989, pages 57–60, Akademie–Verlag, Berlin, 1990.

[63] H.-M. Voigt, H. Mühlenbein, and H.-P. Schwefel, editors. Evolution and Optimization ’89 – Selected Papers on Evolution Theory, Combinatorial Optimization, and Related Topics, Wartburg, April 1989, Akademie–Verlag, Berlin, 1990.

[64] Thomas Bäck, Frank Hoffmeister, and Hans-Paul Schwefel. A survey of evolution strategies. In R. K. Belew and L. B. Booker, editors, Proc. Fourth Int’l Conf. Genetic Algorithms (ICGA’91), San Diego CA, pages 2–9, Morgan Kaufmann, San Mateo CA, 1991.

[65] Hans-Paul Schwefel. Understanding evolution as a collective strategy for groping in the dark. In J. D. Becker, I. Eisele, and F. W. Mündemann, editors, Parallelism, Learning, Evolution, volume 565 of Lecture Notes in Artificial Intelligence, pages 388–397. Springer, Berlin, 1991.

[66] Hans-Paul Schwefel. A toolbox for dynamic modeling. In P. Kleinschmidt and F. J. Radermacher, editors, Proc. Twelfth Symp. Operations Research (SOR), Passau, September 1987, volume 2, pages 479–483, Athenäum, Frankfurt a.M., 1991.

[67] H.-P. Schwefel and R. Männer, editors. Parallel Problem Solving from Nature – Proc. First Workshop PPSN, Dortmund, October 1990, volume 496 of Lecture Notes in Computer Science, Springer, Berlin, 1991.

[68] Thomas Bäck, Frank Hoffmeister, and Hans-Paul Schwefel. Applications of evolutionary algorithms. Technical Report of the Systems Analysis Research Group SYS–2/92, University of Dortmund, Department of Computer Science, February 1992.

[69] Thomas Bäck, Frank Hoffmeister, and Hans-Paul Schwefel. Applications of evolutionary algorithms – 2nd extended edition. Technical Report of the Systems Analysis Research Group SYS–2/92, University of Dortmund, Department of Computer Science, July 1992.

[70] Thomas Bäck and Hans-Paul Schwefel. Evolutionary algorithms: Some very old strategies for optimization and adaptation. In D. Perret-Gallix, editor, New Computing Techniques in Physics Research II – Proc. Second Int’l Workshop Software Engineering, Artificial Intelligence, and Expert Systems for High Energy and Nuclear Physics, pages 247–254, World Scientific, Singapore, 1992.

[71] Yuval Davidor and Hans-Paul Schwefel. An introduction to adaptive optimization algorithms based on principles of natural evolution. In B. Souček and The IRIS Group, editors, Dynamic, Genetic, and Chaotic Programming, pages 183–202. Wiley Interscience, New York, 1992.

[72] Pierre Frankhauser and Hans-Paul Schwefel. Making use of the Weidlich–Haag–model in the case of reduced data sets. In P. Gritzmann, R. Hettich, R. Horst, and E. Sachs, editors, Operations Research’91, 16th Symp., Extended Abstracts, pages 320–323, Physica–Verlag, Heidelberg, 1992.

[73] Hans-Paul Schwefel. Poisk optimuma putyem imitatsii iskusstvennowo intellekta (in Russian). In V. P. Bulatov, editor, Optimizatsiya: Modeli, Metody, Resheniya, Proc. Int’l School ‘Optimization Methods and Their Applications’, Irkutsk, 1989, pages 351–358, Nauka Publ., Novosibirsk, 1992.

[74] Hans-Paul Schwefel. Imitating evolution: Collective, two–level learning processes. In U. Witt, editor, Explaining Process and Change – Approaches to Evolutionary Economics, pages 49–63. The University of Michigan Press, Ann Arbor MI, 1992.

[75] Hans-Paul Schwefel. Natural evolution and collective optimum seeking. In A. Sydow, editor, Computational Systems Analysis – Topics and Trends, pages 5–14. Elsevier, Amsterdam, 1992.

[76] Hans-Paul Schwefel and Thomas Bäck. Künstliche Evolution – eine intelligente Problemlösungsstrategie? KI – Künstliche Intelligenz, 6(2):20–27, 1992.

[77] Hans-Paul Schwefel and Frank Kursawe. Künstliche Evolution als Modell für natürliche Intelligenz. In W. Nachtigall, editor, Technische Biologie und Bionik 1, Proc. 1st Bionik–Congress, pages 73–91, Gustav Fischer, Stuttgart, 1992.

[78] Joachim Sprave and Hans-Paul Schwefel. Evolutionäre Algorithmen auf Transputerfarmen zur Lösung schwieriger Optimierungsprobleme. In R. Grebe and M. Baumann, editors, TAT’92, Abstraktband des 4. bundesweiten Transputer–Anwender–Treffens, pages 106–109. Medical School at the Aachen University of Technology (RWTH), Institute of Physiology, 1992.

[79] Alexander Antamoshkin, Hans-Paul Schwefel, Aimo Törn, George Yin, and Antanas Žilinskas. Systems Analysis, Design, and Optimization – An Introduction. Russian Engineering Academy, Siberian Department, Krasnoyarsk Space Technology University, 1993.

[80] Thomas Bäck, Ulrich Hammel, and Hans-Paul Schwefel. Modelloptimierung mit evolutionären Algorithmen. In A. Sydow, T. Lux, and R. P. Schäfer, editors, Fortschritte in der Simulationstechnik, vol. 6, pages 49–57, Vieweg, Braunschweig, 1993.

[81] Thomas Bäck, Frank Hoffmeister, and Hans-Paul Schwefel. Applications of evolutionary algorithms – 3rd extended edition. Technical Report of the Systems Analysis Research Group SYS–2/92, University of Dortmund, Department of Computer Science, January 1993.

[82] Thomas Bäck, Frank Hoffmeister, and Hans-Paul Schwefel. Applications of evolutionary algorithms – 4th extended edition. Technical Report of the Systems Analysis Research Group SYS–2/92, University of Dortmund, Department of Computer Science, July 1993.

[83] Thomas Bäck, Günter Rudolph, and Hans-Paul Schwefel. Evolutionary programming and evolution strategies: Similarities and differences. In D. B. Fogel and W. Atmar, editors, Proc. Second Annual Conf. Evolutionary Programming (EP’93), pages 11–22, Evolutionary Programming Society, San Diego CA, 1993.

[84] Thomas Bäck and Hans-Paul Schwefel. An overview of evolutionary algorithms for parameter optimization. Evolutionary Computation, 1(1):1–23, 1993.

[85] Hans-Paul Schwefel. Erde – Mensch – Zukunft. In M. Heise and H. Boës, editors, Gedankenfragmente, page 17. Netzwerk Zukunft e.V., Gelsenkirchen, 1993.

[86] Hans-Paul Schwefel and Günter Rudolph. Endbericht der Arbeitsgruppe Schwefel. In P. Drews and B. Monien, editors, Berichte des Forschungsverbundes Nordrhein–Westfalen ‘Paralleles Rechnen’, pages 67–74, Aachen, 1993.

[87] Y. Davidor, H.-P. Schwefel, and R. Männer, editors. Parallel Problem Solving from Nature – PPSN III, Proc. Int’l Conf. Evolutionary Computation, Jerusalem, October 1994, volume 866 of Lecture Notes in Computer Science, Springer, Berlin, 1994.

[88] Ulrich Hammel, Thomas Bäck, and Hans-Paul Schwefel. Evolutionäre Algorithmen: Optimieren nach dem Vorbild der biologischen Evolution. Technical Report of the Systems Analysis Research Group SYS–1/94, University of Dortmund, Department of Computer Science, April 1994.

[89] Z. Michalewicz, J. D. Schaffer, H.-P. Schwefel, D. B. Fogel, and H. Kitano, editors. Proc. First IEEE Conf. Evolutionary Computation (ICEC’94) within First IEEE World Congress on Computational Intelligence (WCCI’94), Orlando FL, June 1994, volume I (oral presentations) and II (posters), IEEE Press, Piscataway NJ, 1994.

[90] Günter Rudolph and Hans-Paul Schwefel. Evolutionäre Algorithmen: Ein robustes Optimierkonzept. Physikalische Blätter, 50(3):236–238, 1994.

[91] Hans-Paul Schwefel. On the evolution of evolutionary computation. In J. M. Zurada, R. J. Marks II, and C. J. Robinson, editors, Computational Intelligence – Imitating Life, pages 116–124. IEEE Press, Piscataway NJ, 1994.

[92] Hans-Paul Schwefel. Inverse Bioinformatik. In F. Haubensak and J. Sühnel, editors, Bioinformatik, Tagungsband: 9. Workshop ‘Computereinsatz in den Biowissenschaften’ und 2. Fachtagung ‘Informatik in den Biowissenschaften’, page 15, Jena, 1994.

[93] Hans-Paul Schwefel, Ulrich Hammel, and Thomas Bäck. Evolutionäre Algorithmen: Optimieren nach dem Vorbild der biologischen Evolution. GMD–Spiegel, 24(1/2):49–58, 1994.

[94] Hans-Paul Schwefel, Ulrich Hammel, and Thomas Bäck. Evolutionäre Algorithmen: Optimieren nach dem Vorbild der biologischen Evolution. In GME–Fachbericht No. 14, Anwendungsfelder in der Mikroelektronik, pages 23–47, VDE/VDI-Gesellschaft Mikroelektronik, Munich, 1994.

[95] George Yin, Günter Rudolph, and Hans-Paul Schwefel. Exploiting the connections of evolutionary algorithms and stochastic approximation. Technical Report of the Systems Analysis Research Group SYS–3/94, University of Dortmund, Department of Computer Science, August 1994.

[96] George Yin, Günter Rudolph, and Hans-Paul Schwefel. Analyzing (1, λ)-evolution strategy via stochastic approximation methods. Technical Report of the Systems Analysis Research Group SYS–4/94, University of Dortmund, Department of Computer Science, August 1994.

[97] Thomas Bäck and Hans-Paul Schwefel. Evolution strategies I: Variants and their computational implementation. In G. Winter, J. Périaux, M. Galán, and P. Cuesta, editors, Genetic Algorithms in Engineering and Computer Science, pages 111–126, Wiley, New York, 1995.

[98] Hans-Paul Schwefel. Evolution and Optimum Seeking. Sixth-Generation Computer Technology. Wiley Interscience, New York, 1995.

[99] Hans-Paul Schwefel et al. Verbundprojekt EVOALG – Grundlagen und Anwendungen evolutionärer Algorithmen. In G. Wolf, R. Schmidt, and M. van der Meer, editors, Statusseminar des BMBF ‘Bioinformatik’, pages 259–280, Projektträger Informationstechnik des BMBF bei der Deutschen Forschungsanstalt für Luft- und Raumfahrt e.V., Berlin, 1995.

[100] Hans-Paul Schwefel and Thomas Bäck. Evolution strategies II: Theoretical aspects. In G. Winter, J. Périaux, M. Galán, and P. Cuesta, editors, Genetic Algorithms in Engineering and Computer Science, pages 127–140, Wiley, New York, 1995.

[101] Hans-Paul Schwefel and Günter Rudolph. Contemporary evolution strategies. In F. Morán, A. Moreno, J. J. Merelo, and P. Chacón, editors, Advances in Artificial Life – Proc. Third European Conf. Artificial Life (ECAL’95), pages 893–907, Springer, Berlin, 1995.

[102] Hans-Paul Schwefel, Günter Rudolph, and Thomas Bäck. Contemporary evolution strategies. Technical Report of the Systems Analysis Research Group SYS–6/95, University of Dortmund, Department of Computer Science, December 1995.

[103] George Yin, Günter Rudolph, and Hans-Paul Schwefel. Establishing connections between evolutionary algorithms and stochastic approximation. Informatica (Lithuania), 6(1):93–117, 1995.

[104] George Yin, Günter Rudolph, and Hans-Paul Schwefel. Analyzing the (1, λ) evolution strategy via stochastic approximation methods. Evolutionary Computation, 3(4):473–489, 1995.

[105] A. Antamoshkin, M. Volovik, V. Ganzhenko, A. Zhilinskas, G. Yin, S. Korobeynikov, A. Torn, H.-P. Schwefel, and V. Sheludko. Sistemy analiz, vol. I and II, Proyektirovaniye, optimizatsiya i prilozheniya. Sibirskaya Aerokosmicheskaya Akademiya, Krasnoyarsk, 1996.

[106] Thomas Bäck, Ulrich Hammel, Martin Schütz, Hans-Paul Schwefel, and Joachim Sprave. Applications of evolutionary algorithms at the Center for Applied Systems Analysis. In J.-A. Désidéri, C. Hirsch, P. Le Tallec, E. Oñate, M. Pandolfi, J. Périaux, and E. Stein, editors, Proc. European Congress on Computational Methods in Applied Sciences: Third Computational Fluid Dynamics Conf. and the Second Conf. Numerical Methods in Engineering (ECCOMAS’96), pages 243–250, Wiley, Chichester, 1996.

[107] Thomas Bäck and Hans-Paul Schwefel. Evolutionary computation: An overview. In T. Fukuda, T. Furuhashi, and D. B. Fogel, editors, Proc. 1996 IEEE Int’l Conf. Evolutionary Computation (ICEC’96), Nagoya, pages 20–29, IEEE Press, Piscataway NJ, 1996.

[108] V. Claus, J. Hopf, and H.-P. Schwefel, editors. Evolutionary Algorithms and Their Application, Dagstuhl-Seminar, March 1996, Report No. 140, IBFI GmbH, Schloss Dagstuhl, Wadern, 1996.

[109] Ulrich Hammel, Thomas Bäck, and Hans-Paul Schwefel. Evolutionary algorithms – Optimization by mimicking biological evolution. GMD–Spiegel: Special Issue on Simulation Models, pages 45–54, 1996.

[110] H.-M. Voigt, W. Ebeling, I. Rechenberg, and H.-P. Schwefel, editors. Parallel Problem Solving from Nature – PPSN IV, Proc. Int’l Conf. Evolutionary Computation, Berlin, September 1996, volume 1141 of Lecture Notes in Computer Science, Springer, Berlin, 1996.

[111] Thomas Bäck, Ulrich Hammel, and Hans-Paul Schwefel. Evolutionary computation: Comments on the history and current state. IEEE Transactions on Evolutionary Computation, 1(1):3–17, 1997.

[112] Kenneth De Jong, David B. Fogel, and Hans-Paul Schwefel. A history of evolutionary computation. In Th. Bäck, D. B. Fogel, and Z. Michalewicz, editors, Handbook of Evolutionary Computation, pages A2.3:1–12. Oxford University Press, New York, and Institute of Physics Publ., Bristol, 1997.

[113] Frank Kursawe and Hans-Paul Schwefel. Optimierung mit evolutionären Algorithmen. atp (Automatisierungstechnische Praxis), 39(9):10–17, 1997.

[114] Ahmet Irfan Oyman, Hans-Georg Beyer, and Hans-Paul Schwefel. Analysis of a simple ES on the parabolic ridge. Technical Report of the Systems Analysis Research Group SYS–2/97, University of Dortmund, Department of Computer Science, August 1997.

[115] Hans-Paul Schwefel. Advantages (and disadvantages) of evolutionary computation over other approaches. In Th. Bäck, D. B. Fogel, and Z. Michalewicz, editors, Handbook of Evolutionary Computation, pages A1.3:1–2. Oxford University Press, New York, and Institute of Physics Publ., Bristol, 1997.

[116] Hans-Paul Schwefel. Challenges to and future developments of evolutionary algorithms. In Th. Bäck, D. B. Fogel, and Z. Michalewicz, editors, Handbook of Evolutionary Computation, pages H1.3:1–4. Oxford University Press, New York, and Institute of Physics Publ., Bristol, 1997.

[117] Hans-Paul Schwefel. Evolutionary computation – A study on collective learning. In N. Callaos, C. M. Khoong, and E. Cohen, editors, Proc. World Multiconference on Systemics, Cybernetics and Informatics (SCI’97), Caracas, volume 2, pages 198–205, Int’l Inst. of Informatics and Systemics (IIIS), Orlando FL, 1997.

[118] Hans-Paul Schwefel. Parallel problem solving from nature. In A. Kent and J. G. Williams, editors, Encyclopedia of Computer Science and Technology, volume 37, supplement 22, pages 225–246. Marcel Dekker, New York, 1997.

[119] Hans-Paul Schwefel. Special track on computational intelligence – Genetic algorithms. In B. Werner, editor, Proc. 23rd Euromicro Conf. – New Frontiers of Information Technology, Budapest, IEEE Computer Society, Los Alamitos CA, 1997.

[120] Hans-Paul Schwefel. Evolutionary computation – A study on collective learning. Technical Report of the Collaborative Research Centre 531 Computational Intelligence CI–13/97, University of Dortmund, November 1997.

[121] Hans-Paul Schwefel and Ulrich Hammel. SFB 531 – Design und Management komplexer Prozesse und Systeme mit Methoden der Computational Intelligence. In M. Jarke, K. Pasedach, and K. Pohl, editors, Informatik’97 – Informatik als Innovationsmotor – 27. Jahrestagung der GI, pages 102–110, Springer, Berlin, 1997.

[122] Hans-Paul Schwefel, Joachim Sprave, Mike Preuß, Susanne Rolf, Rudolf Lewandowski, and Stefan Theis. Projekt PARPROG – Parallele Prognoseverfahren für die Absatzplanung. In R. Krahl and G. Wolf, editors, Statustagung des BMBF: HPSC’97 – Paralleles Höchstleistungsrechnen und seine Anwendungen, pages 211–223, Projektträger des BMBF für Informationstechnik bei der DLR e.V., Berlin, 1997.

[123] Klaus Weinert, Jörn Mehnen, Jens Jasper, Hans-Paul Schwefel, and Hans-Georg Beyer. Entwicklung von Prozessmodellen der spanenden Bearbeitung mit Methoden der nichtlinearen Dynamik. In 1. Symposium des Förderschwerpunktes ‘Untersuchung nichtlinear–dynamischer Effekte in produktionstechnischen Systemen’ der Volkswagen Stiftung, pages 1–24. University of Hanover, Institute of Production Systems, 1997.

[124] A. E. Eiben, Th. Bäck, M. Schoenauer, and H.-P. Schwefel, editors. Parallel Problem Solving from Nature – PPSN V, Proc. Fifth Int’l Conf., Amsterdam, September 1998, volume 1498 of Lecture Notes in Computer Science, Springer, Berlin, 1998.

[125] D. B. Fogel, H.-P. Schwefel, Th. Bäck, and X. Yao, editors. Proc. Fifth IEEE Conf. Evolutionary Computation (ICEC’98) within Second IEEE World Congress on Computational Intelligence (WCCI’98), Anchorage AK, May 1998, IEEE Press, Piscataway NJ, 1998.

[126] Sigrid Hafner, Harro Kiendl, and Hans-Paul Schwefel. Computational Intelligence: Neuronale Netze, Evolutionäre Algorithmen und Fuzzy Control wachsen zusammen. atp (Automatisierungstechnische Praxis), 40(2):8–10, 1998.

[127] S. Hafner, H. Kiendl, and H.-P. Schwefel, editors. Computational Intelligence: Neuronale Netze, Evolutionäre Algorithmen, Fuzzy Control im industriellen Einsatz, Proc. VDI-VDE-GMA-Fachtagung, Berlin, March 1998, volume 1381 of VDI-Berichte, VDI-Verlag, Düsseldorf, 1998.

[128] Frank Kursawe and Hans-Paul Schwefel. Künstliche Evolution als Modell für natürliche Intelligenz. In A. v. Gleich, editor, Bionik – Ökologische Technik nach dem Vorbild der Natur?, pages 65–89. Teubner, Stuttgart, 1998.

[129] Frank Kursawe and Hans-Paul Schwefel. Optimierung mit evolutionären Algorithmen. In S. Hafner, editor, Industrielle Anwendungen evolutionärer Algorithmen, pages 1–13. R. Oldenbourg, München, 1998.

[130] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. A spatial predator-prey approach to multi-objective optimization. In A. E. Eiben, Th. Bäck, M. Schoenauer, and H.-P. Schwefel, editors, Proc. Parallel Problem Solving from Nature – PPSN V, Amsterdam, pages 241–249, Springer, Berlin, 1998.

[131] Ahmet Irfan Oyman, Hans-Georg Beyer, and Hans-Paul Schwefel. Convergence behavior of the (1 ,^{+} λ)-evolution strategy on the ridge functions. Technical Report of the Systems Analysis Research Group SYS–1/98, University of Dortmund, Department of Computer Science, February 1998.

[132] Ahmet Irfan Oyman, Hans-Georg Beyer, and Hans-Paul Schwefel. Where elitists start limping: Evolution strategies at ridge functions. In A. E. Eiben, Th. Bäck, M. Schoenauer, and H.-P. Schwefel, editors, Proc. Parallel Problem Solving from Nature – PPSN V, Amsterdam, pages 34– 43, Springer, Berlin, 1998.

[133] Hans-Paul Schwefel. Evolutionäres Rechnen in der Computational Intelligence. In S. Hafner, H. Kiendl, and H.-P. Schwefel, editors, Computational Intelligence: Neuronale Netze, Evolutionäre Algorithmen, Fuzzy Control im industriellen Einsatz, pages 19–23, VDI-Verlag, Düsseldorf, 1998.

[134] Hans-Paul Schwefel and Thomas Bäck. Artificial evolution: How and why? In D. Quagliarella, J. Périaux, C. Poloni, and G. Winter, editors, Genetic Algorithms and Evolution Strategies in Engineering and Computer Science, pages 1–19, Wiley, Chichester, 1998.

[135] Hans-Paul Schwefel and Ulrich Hammel. Design und Management komplexer Prozesse und Systeme mit Methoden der Computational Intelligence. Technical Report of the Collaborative Research Centre 531 Computational Intelligence CI–26/98, University of Dortmund, April 1998.

[136] Hans-Paul Schwefel and Frank Kursawe. On natural life’s tricks to survive and evolve. Technical Report of the Collaborative Research Centre 531 Computational Intelligence CI–18/98, University of Dortmund, January 1998.

[137] Hans-Paul Schwefel and Frank Kursawe. On natural life’s tricks to survive and evolve. In D. B. Fogel, H.-P. Schwefel, Th. Bäck, and X. Yao, editors, Proc. Fifth IEEE Conf. Evolutionary Computation (ICEC’98), Anchorage AK, pages 1–8, IEEE Press, Piscataway NJ, 1998.

[138] Klaus Weinert, Jörn Mehnen, Martin Kalveram, Hans-Paul Schwefel, and Boris Naujoks. Analyse des HSC-Fräsprozesses mit Hilfe der Methoden der Nichtlinearen Dynamik. VDI-Z Integrierte Produktion, 140(9):54–57, 1998.

[139] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. Approximating the Pareto set: Concepts, diversity issues, and performance assessment. Technical Report of the Collaborative Research Centre 531 Computational Intelligence CI–72/99, University of Dortmund, March 1999.

[140] Hans-Paul Schwefel. Natural computation (abstract). In B. McKay, X. Yao, C. S. Newton, J.-H. Kim, and T. Furuhashi, editors, Simulated Evolution and Learning; selected papers – Second Asia-Pacific Conf. (SEAL’98), volume 1585 of Lecture Notes in Computer Science, page 1, Springer, Berlin, 1999.

[141] Hans-Paul Schwefel. Introduction to evolution strategies. In Proc. Genetic and Evolutionary Computation Conf. (GECCO 1999) – Tutorial Program, pages 321–346, Orlando FL, 1999.

[142] Hans-Paul Schwefel. Evolving solutions for design and management tasks on computers. In Proc. IEEE Int’l Conf. Systems, Man, and Cybernetics (SMC’99), volume III, pages 573–578, IEEE, Tokyo, 1999. (CD-ROM Catalog Number 99CH37028C).

[143] Hans-Paul Schwefel. Evolving solutions for design and management tasks on computers. Technical Report of the Collaborative Research Centre 531 Computational Intelligence CI–71/99, University of Dortmund, July 1999.

[144] Hans-Paul Schwefel, Joachim Sprave, Rudolf Lewandowski, and Stefan Theis. Parallele Prognoseverfahren für die Absatzplanung – Schlussbericht. Technical Report of the Systems Analysis Research Group SYS–2/99, University of Dortmund, Department of Computer Science, July 1999.

[145] Klaus Weinert, Jörn Mehnen, Martin Kalveram, Hans-Paul Schwefel, and Boris Naujoks. Nonlinear time series analysis of milling processes. In F. Klocke, editor, Second Int’l Symp. Investigation of Nonlinear Dynamic Effects in Production Systems, chapter 6. Laboratory of Machining Tools and Production Engineering (WZL), Aachen University of Technology (RWTH), 1999.

[146] Klaus Weinert, Jörn Mehnen, Martin Kalveram, Hans-Paul Schwefel, and Boris Naujoks. On the analysis of the dynamics of high-speed milling by means of time series analysis methods. In Production Engineering – Research and Development, Annals of the German Academic Society for Production Engineering, volume VI/2, pages 21–24. Wissenschaftliche Gesellschaft für Produktionstechnik, Berlin, 1999.

[147] Kenneth De Jong, David B. Fogel, and Hans-Paul Schwefel. A history of evolutionary computation. In Th. Bäck, D. B. Fogel, and Z. Michalewicz, editors, Evolutionary Computation 1 – Basic Algorithms and Operators, pages 40–58. Institute of Physics Publ., Bristol, 2000.

[148] S. Hafner, H. Kiendl, R. Kruse, and H.-P. Schwefel, editors. Computational Intelligence im industriellen Einsatz: Fuzzy Systeme, Neuronale Netze, Evolutionäre Algorithmen, Data Mining, Proc. VDI-VDE-GMA-Fachtagung, Baden-Baden, May 2000, volume 1526 of VDI-Berichte, VDI-Verlag, Düsseldorf, 2000.

[149] Ahmet Irfan Oyman, Hans-Georg Beyer, and Hans-Paul Schwefel. Analysis of the (1, λ)-ES on the parabolic ridge. Evolutionary Computation, 8(3):249–265, 2000.

[150] M. Schoenauer, K. Deb, G. Rudolph, X. Yao, E. Lutton, J. J. Merelo, and H.-P. Schwefel, editors. Parallel Problem Solving from Nature – PPSN VI, Proc. Sixth Int’l Conf., Paris, September 2000, volume 1917 of Lecture Notes in Computer Science, Springer, Berlin, 2000.

[151] Hans-Paul Schwefel. Advantages (and disadvantages) of evolutionary computation over other approaches. In Th. Bäck, D. B. Fogel, and Z. Michalewicz, editors, Evolutionary Computation 1 – Basic Algorithms and Operators, pages 20–22. Institute of Physics Publ., Bristol, 2000.

[152] Martin Schütz and Hans-Paul Schwefel. Evolutionary approaches to solve three challenging engineering tasks. CMAME (Computer Methods in Applied Mechanics and Engineering), 186:141–170, 2000.

[153] Frank Kursawe and Hans-Paul Schwefel. Künstliche Evolution als Modell für natürliche Intelligenz. In A. v. Gleich, editor, Bionik – Ökologische Technik nach dem Vorbild der Natur?, pages 65–85. Teubner, Stuttgart, 2nd edition, 2001.

[154] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. Mutation control and convergence in evolutionary multi-objective optimization. In R. Matousek and P. Osmera, editors, Proc. Seventh Int’l Conf. Soft Computing (MENDEL’01), pages 24–29, University of Technology, Brno, Czech Republic, 2001.

[155] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. Mutation control and convergence in evolutionary multi-objective optimization. Technical Report of the Collaborative Research Centre 531 Computational Intelligence CI–109/01, University of Dortmund, April 2001.

[156] Hans-Georg Beyer and Hans-Paul Schwefel. Evolution strategies – A comprehensive introduction. Natural Computing, 1:3–52, 2002.

[157] Hans-Georg Beyer, Hans-Paul Schwefel, and Ingo Wegener. How to analyse evolutionary algorithms. Theoretical Computer Science, 287:101–130, 2002.

[158] Hans-Georg Beyer, Hans-Paul Schwefel, and Ingo Wegener. How to analyse evolutionary algorithms. Technical Report of the Collaborative Research Centre 531 Computational Intelligence CI–139/02, University of Dortmund, August 2002.

[159] J. J. Merelo Guervós, P. Adamidis, H.-G. Beyer, J. L. Fernández-Villacañas, and H.-P. Schwefel, editors. Parallel Problem Solving from Nature – PPSN VII, Proc. Seventh Int’l Conf., Granada, volume 2439 of Lecture Notes in Computer Science, Springer, Berlin, 2002.

[160] Hans-Paul Schwefel. Deep insight from simple models of evolution. BioSystems, 64(1–3):189–198, [161] Thomas Beielstein, Jörn Mehnen, Lutz Schönemann, Hans-Paul Schwefel, Tobias Surmann, Klaus

Weinert, and Dirk Wiesmann. Design of evolutionary algorithms and applications in surface reconstruction. In H.-P. Schwefel, I. Wegener, and K. Weinert, editors, Advances in Computational Intelligence – Theory and Practice, pages 145–193. Springer, Berlin, 2003.

[162] Stefan Droste, Thomas Jansen, Günter Rudolph, Hans-Paul Schwefel, Karsten Tinnefeld, and Ingo Wegener. Theory of evolutionary algorithms and genetic programming. In H.-P. Schwefel, I. Wegener, and K. Weinert, editors, Advances in Computational Intelligence – Theory and Practice, pages 107–144. Springer, Berlin, 2003.

[163] Ulrich Hammel, Boris Naujoks, and Hans-Paul Schwefel. Introduction. In H.-P. Schwefel, I. Wegener, and K. Weinert, editors, Advances in Computational Intelligence – Theory and Practice, pages 1–11. Springer, Berlin, 2003.

[164] H.-P. Schwefel, I. Wegener, and K. Weinert, editors. Advances in Computational Intelligence – Theory and Practice. Natural Computing Series. Springer, Berlin, 2003.

[165] X. Yao, E. Burke, J. A. Lozano, J. Smith, J. J. Merelo Guervós, J. A. Bullinaria, J. Rowe, P. Tiňo, A. Kabán, and H.-P. Schwefel, editors. Parallel Problem Solving from Nature – PPSN VIII, Proc. Eighth Int’l Conf., Birmingham, volume 3242 of Lecture Notes in Computer Science, Springer, Berlin, 2004.

[166] Oliver Kramer and Hans-Paul Schwefel. On three new approaches to handle constraints within evolution strategies. Natural Computing, 5:363–385, 2006.

[167] Günter Rudolph and Hans-Paul Schwefel. Simulated evolution under multiple criteria conditions revisited. In J. M. Zurada, G. G. Yen, and J. Wang, editors, Computational Intelligence – Research Frontiers, Plenary and Invited Lectures at the IEEE World Congress on Computational Intelligence, WCCI 2008, Hong Kong, June 1–6, 2008, volume 5050 of Lecture Notes in Computer Science, State-of-the-Art-Survey, pages 249–261. Springer, Berlin, 2008.

[168] Hans-Paul Schwefel. Foreword. In Muddassar Farooq, editor, Bee-Inspired Protocol Engineering – From Nature to Networks, Natural Computing Series, page VII. Springer, Berlin, 2009.

[169] Hans-Paul Schwefel. eVolution seit dem Beginn der Informatik – Abstract. In S. Fischer, E. Maehle, and R. Reischuk, editors, Informatik 2009 – Im Focus das Leben – Proceedings of the 39th Annual Conference of the GI, Lubeck, 28 Sept.–2 Oct. 2009, Lecture Notes in Informatics (LNI), page 59, Gesellschaft für Informatik e.V. (GI), Bonn, 2009.

[170] Hans-Paul Schwefel and Uwe Schwiegelshohn. Editorial Introduction to the Special Issue: Evolutionary Computing in the Collaborative Research Centre on Computational Intelligence at Technische Universität (TU) Dortmund. Evolutionary Computation, 17(4):iii–iv, 2009.

[171] Thomas Bartz-Beielstein, Mike Preuß, Karlheinz Schmitt, and Hans-Paul Schwefel. Model Optimization with Evolutionary Algorithms. In Klaus Lucas and Peter Roosen, editors, Emergence, Analysis, and Evolution of Structures – Concepts and Strategies Across Disciplines, Springer Complexity Program, pages 47–62. Springer, Heidelberg, 2010.

[172] Thomas Bartz-Beielstein, Mike Preuss, Karlheinz Schmitt, and Hans-Paul Schwefel. Challenges for Contemporary Evolutionary Algorithms. Technical Report of Chair Algorithm Engineering (Ls11) TR10-2-003, Dortmund University of Technology, Faculty of Computer Science, May 2010.

[173] Hans-Paul Schwefel. Foreword. In Thomas Bartz-Beielstein, Marco Chiarandini, Luís Paquete, and Mike Preuss, editors, Experimental Methods for the Analysis of Optimization Algorithms, pages vii– viii. Springer, Berlin, 2010.

[174] Hans-Paul Schwefel. Darwin auf dem Computer: Evolutionäre Algorithmen. In Volker Gerhardt, Klaus Lucas, and Günter Stock, editors, Evolution–Theorie, Formen und Konsequenzen eines Paradigmas in Natur, Technik und Kultur, pages 119–132 and 293–294. Akademie Verlag of the berlin-brandenburgische Akademie der Wissenschaften (BBAW), Berlin, 2011.

[175] Hans-Paul Schwefel. Life lessons taught by simulated evolution. In Mark Burgin and Eugene Eberbach, editors, Evolutionary Computation and the Processes of Life, pages 1–9. Association for Computing Machinery (ACM) Ubiquity Symposium, New York NY, September 2012.

[176] Hans-Paul Schwefel. Foreword. In Pierre Collet and Shigeyoshi Tsutsui, editors, Massively Parallel Evolutionary Computation on GPGPUs, pages vii–viii. Springer, Berlin, 2013.

[177] Hans-Paul Schwefel. Foreword. In Mike Preuss, editor, Multimodal Optimization by Means of Evolutionary Algorithms, page VII. Springer, Int’l, 2015.

[178] Hans-Paul  Schwefel.  Bionik,  Kybernetik

und Evolutionäre Algorithmen. In M. Lehmann-Waffenschmidt and M. Paneder, editors, Evolutorische Ökonomik, pages 105–114. Springer Gabler, Wiesbaden, 2022. (eBook 978-3-658-34287-6)

Publications about or for Hans–Paul Schwefel.

[179] Karl Steinbuch. Die informierte Gesellschaft. pages 277–282. Deutsche Verlags-Anstalt DVA, Stuttgart, 1966.

[180] Karl Steinbuch. Die informierte Gesellschaft. pages 202–205. DVA rororo Taschenbuch Ausgabe, Reinbek, 1968.

[181] Thomas Bartz-Beielstein, Gundel Jankord, Boris Naujoks, Günter Rudolph, and Karlheinz Schmitt, editors. Festschrift Hans–Paul Schwefel 2006. SFB 531 Computational Intelligence. University of Dortmund, Chair of Systems Analysis, 2006.

[182] Pier Luca Lanzi, Günter Rudolph, and Hans-Paul Schwefel. An Interview with Hans–Paul Schwefel. SIGEVOlution, Newsletter of the ACM Special Interest Group on Genetic and Evolutionary Computation, 3(4):2–5, 2008.

[183] Marcelo Augusto De Brito Mendes. Forty Five Years Of Evolution Strategies. Blogspot from Brazil. Genetic Argonaut, http://geneticargonaut.blogspot.com/, June 12, 2009.

[184] Hans-Paul Schwefel. Lebenslauf von eigener Hand. http://hdl.handle.net/2003/26415. Biografisches Archiv Dortmunder Universitätsprofessorinnen und -professoren, Eldorado – Ressourcen aus und für Lehre, Studium und Forschung, Technische Universität Dortmund, September 15, 2009.

[185] Marcelo Augusto De Brito Mendes and Hans-Paul Schwefel. 45 Years of Evolution Strategies. SIGEVOlution, Newsletter of the ACM Special Interest Group on Genetic and Evolutionary Computation, 4(2):2–8, 2009.